# ChemGen: Code Generation for Multispecies Chemical Kinetics in Physics Simulations

### v1

@@ 6 Conclusion
We validated the correctness of the generated code by comparing its results against established Cantera source terms and showed consistency across a range of reaction mechanisms commonly used in CFD. We further verified the accuracy of the Jacobian computations and demonstrated optimal convergence across several supplied implicit integration methods. Finally, we embedded ChemGen into an established CFD code to perform a detonation simulation and reproduced results consistent with those in the literature, illustrating its ability to integrate seamlessly into mature computational frameworks. All data, source code, and supplementary tutorials are publicly available.
ChemGen lays the groundwork for the consistent inclusion of chemical source terms in diverse computational environments. Its flexibility also opens the door for scalable solvers, reduced-order modeling strategies, and novel chemistry integration methods that can advance the frontier of high-fidelity chemical reacting flow simulation.
@@ A.1 SDIRK-2
The two stage Singly Diagonally Implicit Runge–Kutta (SDIRK-2) method belongs to a class of implicit Runge-Kutta schemes [40, 60]. These methods are stable and exhibit higher order of accuracy than backward Euler, but are more expensive. The SDIRK-2 method is as follows:
@@
Here $\omega_{i}$ is evaluated from the initial guess of $\Gamma_{2,k=1}=\mathcal{S}\left(y_{n}\right)$ with subsequent $\Gamma_{2,k}$ coming from the iterated for decrement, $\Gamma_{2,k+1}=\Gamma_{2,k}+\delta\Gamma_{2}$. Once $\Gamma_{2,k}$ is realized the next stage can be solved for in a similar manner. Chemgen utilizes $\gamma=1-\sqrt{2}/2$ [41].
@@ A.2 SDIRK-4
The five stage Singly Diagonally Implicit Runge-Kutta (SDIRK-4) follows the same structure as SDIRK-2 with additional stages requiring a set of coefficients. We utilize the coefficients implemented by FATODE [41] which have shown stability for stiff problems and achieve 4th order accuracy. The SDIRK-4 method is written as
@@
Here $\mathcal{J}_{n}$ is the chemical Jacobian, $\mathcal{J}_{ij}$,held constant for both stages, thus removing the need for multiple Jacobian evaluations. Again we utilize the coefficients from FATODE [41],
@@ C.1 Source term error due to thermodynamic polynomial degree
Figure 10: Error distributions (Equation (62)) for ChemGen source calculations in comparison to Cantera source calculations for five tested chemical models: Ó Conaire [54], Burke [55], GRI-Mech 3.0 [56], UCSD [57], and FFCM2 [20]. The vertical dash line represents the mean of the distributions and is colored according to the corresponding model. The distribution mean, $\mu$, and corresponding standard deviation, $\sigma$, are also reported.
@@
we note that the very small value of $T_{3}$ effectively drives the first term to zero, while the extremely large value of $T_{1}$ drives the second exponential term to one. It is assumed, based on the specification, that omitting $T_{2}$ entirely implies that the third term is excluded from the calculation, rather than setting $T_{2}=0$ [22]. Updating ChemGen to address these edge cases and avoid numerical instabilities led to reduced error and recovered the expected trend that higher-order polynomial fits improved accuracy.
@@ C.2 Source term Jacobian error
Niemeyer et al. verified the PyJac-generated Jacobian by comparing their analytical Jacobian to a differentiated Jacobian across a wide range of thermochemical states, evaluating between 100,000 and 1 million samples depending on the chemical model [24]. They reported two metrics that could be used per chemical state to evaluate the relative error,
@@
Mechanism $E_{norm}$ Mean Error $E_{norm}$ Max Error $E_{rel}$ Mean Error $E_{rel}$ Max Error

### v3

@@ 6 Conclusion
We validated the correctness of the generated code by comparing its results against established Cantera source terms and showed consistency across a range of reaction models commonly used in CFD. We further verified the accuracy of the Jacobian computations and demonstrated optimal convergence across several supplied implicit integration methods. Finally, we embedded ChemGen into an established CFD code to perform a detonation simulation and reproduced results consistent with those in the literature, illustrating its ability to integrate seamlessly into mature computational frameworks. All data, source code, and supplementary tutorials are publicly available.
ChemGen lays the groundwork for the consistent inclusion of chemical source terms in diverse computational environments. Its flexibility also opens the door for scalable solvers, reduced-order modeling strategies, and novel chemistry integration methods that can advance the frontier of high-fidelity chemical reacting flow simulation. ChemGen is an ongoing project released under the NRL Open License, a source-available license provided by the U.S. Naval Research Laboratory.
7 Acknowledgment
The work at Stanford is sponsored by Office of Naval Research under grants N00014-21-1-2475, N00014-22-1-2606 and N00014-23-1-2501. ESG and JEL also acknowledge support from the National Science Foundation under Grant No. DGE-1656518 and Grant No. DGE-2146755, respectively. RFJ, EJC, and JA acknowledge Dr. Eric Marineau of the Hypersonic Aerothermodynamics, High-Speed Propulsion and Materials Program of the Office of Naval Research Code 35 for support of this work.
@@ A.1 SDIRK-2
The two stage Singly Diagonally Implicit Runge–Kutta (SDIRK-2) method belongs to a class of implicit Runge-Kutta schemes [Hai96, Fran97]. These methods are stable and exhibit higher order of accuracy than backward Euler, but are more expensive. The SDIRK-2 method is as follows:
@@
Here $\omega_{i}$ is evaluated from the initial guess of $\Gamma_{2,k=1}=\mathcal{S}\left(y_{n}\right)$ with subsequent $\Gamma_{2,k}$ coming from the iterated for decrement, $\Gamma_{2,k+1}=\Gamma_{2,k}+\delta\Gamma_{2}$. Once $\Gamma_{2,k}$ is realized the next stage can be solved for in a similar manner. Chemgen utilizes $\gamma=1-\sqrt{2}/2$ [Zha14].
@@ A.2 SDIRK-4
The five stage Singly Diagonally Implicit Runge-Kutta (SDIRK-4) follows the same structure as SDIRK-2 with additional stages requiring a set of coefficients. We utilize the coefficients implemented by FATODE [Zha14] which have shown stability for stiff problems and achieve 4th order accuracy. The SDIRK-4 method is written as
@@
Here $\mathcal{J}_{n}$ is the chemical Jacobian, $\mathcal{J}_{ij}$,held constant for both stages, thus removing the need for multiple Jacobian evaluations. Again we utilize the coefficients from FATODE [Zha14],
@@ C.1 Source term error due to thermodynamic polynomial degree
Figure 10: Error distributions (Equation (63)) for ChemGen source calculations in comparison to Cantera source calculations for five tested chemical models: Ó Conaire [Oco04], Burke [Bur12], GRI-Mech 3.0 [Smi00], UCSD [SanDiegoMech], and FFCM2 [Zha23]. The vertical dash line represents the mean of the distributions and is colored according to the corresponding model. The distribution mean, $\mu$, and corresponding standard deviation, $\sigma$, are also reported.
@@
we note that the very small value of $T_{3}$ effectively drives the first term to zero, while the extremely large value of $T_{1}$ drives the second exponential term to one. It is assumed, based on the specification, that omitting $T_{2}$ entirely implies that the third term is excluded from the calculation, rather than setting $T_{2}=0$ [cantera]. Updating ChemGen to address these edge cases and avoid numerical instabilities led to reduced error and recovered the expected trend that higher-order polynomial fits improved accuracy.
@@ C.2 Source term Jacobian error
Niemeyer et al. verified the PyJac-generated Jacobian by comparing their analytical Jacobian to a differentiated Jacobian across a wide range of thermochemical states, evaluating between 100,000 and 1 million samples depending on the chemical model [Nie17]. They reported two metrics that could be used per chemical state to evaluate the relative error,
@@
Model $E_{norm}$ Mean Error $E_{norm}$ Max Error $E_{rel}$ Mean Error $E_{rel}$ Max Error